# INSIGHT: Explainable Weakly-Supervised Medical Image Analysis

### v2

@@ 4.6 Ablation Studies
To validate the effectiveness of INSIGHT’s architectural components, we perform an ablation study analyzing three key innovations: context suppression, SmoothMax pooling, and regularization. Results for the MosMed, CAMELYON16, and BRACS datasets are shown in Table 3. The study evaluates both binary (MosMed, CAMELYON16) and multi-label (BRACS) tasks, revealing how each component contributes to classification and segmentation performance.
Key Ablation Study Findings:
@@
  • Context Suppression: Adding context suppression (Row 2) improves Dice scores across all datasets, with a notable gain of $16.6$ on CAMELYON16. This enhancement highlights the module’s ability to improve spatial alignment between predictions and ground truth by selectively focusing on relevant regions while suppressing background noise. However, context suppression alone does not maximize AUC, indicating that spatial coherence alone is insufficient for capturing the discriminative features required for optimal classification performance.
@@
  • SmoothMax Pooling: Incorporating SmoothMax pooling alongside context suppression (Row 3) results in a substantial Dice score increase of $34.6$ on CAMELYON16, emphasizing its role in refining lesion boundaries by leveraging all pixel-level activations. This pooling strategy improves the model’s ability to capture nuanced spatial details, enhancing boundary precision. However, SmoothMax pooling slightly reduces AUC for binary tasks, likely due to the softened decision boundaries introduced by spatial averaging, which may blur class distinctions.
@@
  • Regularization: Adding regularization (Row 4) improves both AUC and Dice scores across datasets, stabilizing AUC performance in MosMed and enhancing robustness in BRACS. Spectral decoupling mitigates overfitting in single-label tasks like MosMed, ensuring broader feature utilization and improved generalization. For multi-label tasks like BRACS, label smoothing particularly benefits smaller classes such as ADH and FEA, improving classification consistency across imbalanced distributions.
5 Discussion
@@
Accurately detecting and interpreting disease-related regions in medical imaging remains a critical challenge, particularly when only image-level annotations are available. INSIGHT addresses this gap by achieving state-of-the-art performance in both classification and segmentation tasks across CT volumes and WSIs using only image-level labels. This capability is transformative, as it aligns with real-world clinical workflows where physicians rarely annotate images at the pixel or voxel level. By enabling the use of far larger datasets for training, INSIGHT unlocks the potential for scalable and efficient medical imaging solutions.
INSIGHT’s architectural innovations—its detection and context modules—allow it to combine local feature analysis with contextual understanding, generating interpretable and well-calibrated heatmaps. These heatmaps not only enhance diagnostic transparency but also reduce reliance on extensive pixel-level annotations. Such capabilities could accelerate physician annotations [30], allowing clinicians to correct heatmaps rather than annotate from scratch, thereby streamlining annotation workflows where pixel-level labels are required.
Foundation Models and Future Potential.
For CT scans, we employed DINOv2 pre-trained on ImageNet due to the absence of publicly available radiology-specific pre-trained models. While this approach yielded strong results, we anticipate further improvements with embeddings tailored to radiology data, which could better capture salient features in CT scans. Similarly, for WSIs, we used UNI, one of the first foundation models in pathology. However, newer and more powerful models, such as Virchow [41], have recently emerged. Leveraging these advanced models for feature extraction could significantly enhance INSIGHT’s performance. Unfortunately, due to computational constraints and the recent public release of these models, we were unable to perform a comparative analysis. Future work will explore integrating these cutting-edge foundation models into INSIGHT to fully harness their potential.
Broader Applicability and Limitations.
@@
While INSIGHT has been validated on CT and WSI data, its applicability to other imaging modalities, such as MRI and ultrasound, remains unexplored. In principle, INSIGHT’s architecture should generalize well to these modalities due to its ability to process spatially embedded features and generate interpretable outputs. However, a significant barrier to extending INSIGHT lies in the scarcity of publicly available datasets for these modalities. Future work will focus on identifying or curating appropriate datasets to evaluate INSIGHT’s generalizability and performance across a broader range of imaging modalities.
6 Conclusion
@@
INSIGHT represents a significant advancement in weakly supervised medical imaging, demonstrating robust performance across diverse tasks and modalities. By effectively leveraging image-level labels, integrating interpretability directly into its architecture, and providing fine-grained diagnostic insights, INSIGHT lays the foundation for scalable and clinically relevant AI solutions. Future efforts will aim to enhance its capabilities with advanced foundation models and extend its applicability to additional imaging modalities, furthering its impact on medical imaging research and practice.
Acknowledgments.
@@
Appendix A Qualitative Analysis
@@
We provide additional visualization results (Fig. 345) to further demonstrate INSIGHT’s ability to generate well-calibrated heatmaps that align with diagnostically relevant regions in the Camelyon16 dataset. These visualizations reinforce the interpretability of INSIGHT. Please zoom in to view finer details.
Figure 3: WSI heatmap visualizations generated by INSIGHT. The green outlines on the original image represent ground-truth regions.
@@
Figure 4: Original histopathology image (left) and the corresponding heatmap generated by INSIGHT (right). The green outlines on the original image represent ground-truth regions. The bottom row provides zoomed-in views of the regions highlighted by yellow boxes.
@@
Figure 5: Additional WSI heatmap visualizations generated by INSIGHT. The green outlines represent ground-truth regions.
Appendix B Ablation
INSIGHT retains spatial resolution during the feature extraction process, and the output feature maps preserve spatial structure, resulting in dimensions of $16\times 16\times 1024$ for CT slices and $14\times 14\times 1024$ for WSI tiles. In contrast, compared methods typically apply a pooling strategy that reduces the feature map to a $1\times 1\times 1024$ vector that discards spatial information. To evaluate the effectiveness of our spatial embedding, we conducted an ablation study where the spatial embedding was replaced with a non-spatial one. The results are summarized in Table 4, which highlights the impact of maintaining spatial resolution. Retaining spatial resolution increased the classification AUC by 4.5% and improved the Dice score for segmentation by over 15%. These substantial improvements underscore the critical role of spatial information in histopathological analysis, particularly for segmentation tasks that rely on fine-grained details and spatial continuity to localize regions of interest accurately. Additionally, compared to transformer-based aggregators, which face a quadratic increase in computational cost, INSIGHT leverages its lightweight CNN architecture to achieve a near-linear increase in computation, enabling it to fully capitalize on the benefits of spatial embedding.
Table 4: Ablation study results comparing tensor embeddings and vector embeddings for WSIs on CAMELYON16 dataset. We report AUC for classification, Dice for segmentation, and average time used for generating a heatmap.
Spatial Embedding AUC Dice (%) Runtime (s)
0.945 59.1 $\pm$ 24.8 1
0.990 74.6 $\pm$ 19.1 33

### v3

@@ 5.5 Ablation Studies
To validate the effectiveness of INSIGHT’s architectural components, we perform an ablation study analyzing three key innovations: context suppression, SmoothMax pooling, and regularization. Results for the CAMELYON16 and BRACS datasets are shown in Table 6. The study evaluates both binary (CAMELYON16) and multi-label (BRACS) tasks, revealing how each component contributes to classification and segmentation performance. Our key findings are the following.
Table 5: Comparison of Dice scores (mean $\pm$ std) between INSIGHT’s built-in heatmaps and Grad-CAM. Results are reported on the MosMed and CAMELYON16 datasets.
Dataset Foundation Model Built-in Heatmap Grad-CAM
MosMed DINOv2 42.7 $\pm$ 15.3 12.8 $\pm$ 13.9
Camelyon16 UNI 74.6 $\pm$ 19.1 26.4 $\pm$ 20.4
Camelyon16 Virchow2 78.3 $\pm$ 15.4 20.6 $\pm$ 27.4
Table 6: Ablation study results on the CAMELYON16 and BRACS datasets. INSIGHT performance is shown with different configurations of CS (Context Suppression), SM (SmoothMax), and Rg (Regularizer).
CAMELYON16 BRACS
CS SM Rg AUC Dice(%) ADH FEA DCIS Invasive
[…]
0.982 $\pm21.4$ 0.733 0.507 0.799 0.998
[…]
@@
Context Suppression. Adding context suppression (Row 2) improves Dice scores across all datasets, with a notable gain of $16.6$ on CAMELYON16. This highlights the module’s ability to improve spatial alignment between predictions and ground truth by focusing on relevant regions while suppressing background noise. However, context suppression alone does not maximize AUC, indicating spatial coherence is insufficient for capturing the discriminative features needed for optimal classification.
@@
SmoothMax Pooling. Incorporating SmoothMax pooling alongside context suppression (Row 3) yields a substantial Dice increase of $34.6$ on CAMELYON16, emphasizing its role in refining lesion boundaries by leveraging all pixel-level activations. This pooling strategy enhances the model’s ability to capture nuanced spatial details and boundary precision. However, it slightly reduces AUC for binary tasks, likely due to softened decision boundaries introduced by spatial averaging, which may blur class distinctions. Its ablation (“SmoothMax Removal”) reverts to standard max pooling.
@@
Regularization. Adding regularization (Row 4) improves both AUC and Dice across datasets, stabilizing AUC in CAMELYON16 and enhancing robustness in BRACS. Spectral decoupling mitigates overfitting in single-label tasks like CAMELYON16, ensuring broader feature utilization and improved generalization. For multi-label tasks like BRACS, label smoothing particularly benefits smaller classes such as ADH and FEA, improving classification consistency across imbalanced distributions.
Ablation with Virchow2 Backbone. To further evaluate the generalizability of INSIGHT’s architectural components, we repeat our ablation study using the Virchow2 (Zimmermann et al., 2024) foundational model. This experiment is conducted on CAMELYON16 to assess how our modules perform under a different pretraining regime. As shown in Table 7, we observe consistent performance gains with the addition of each module. The full INSIGHT configuration, incorporating context suppression, SmoothMax pooling, and regularization, achieves the highest overall performance with an AUC of $0.992$ and a Dice score of $78.3\%$. These results reinforce the robustness of our architecture and demonstrate that INSIGHT’s design principles are effective across diverse pretrained backbones.
Table 7: Ablation study of INSIGHT using the Virchow2 foundational model on the Camelyon16 dataset.
Virchow2
CS SM Rg AUC Dice (%)
[…]
6 Discussion and Conclusion
@@
Accurately detecting and interpreting disease-related regions in medical imaging remains a critical challenge, particularly when only image-level annotations are available. INSIGHT addresses this gap by achieving state-of-the-art performance in classification and segmentation across CT volumes and WSIs using only image-level labels. This capability is transformative, aligning with real-world clinical workflows where physicians rarely annotate at the pixel or voxel level. By enabling use of larger datasets, INSIGHT unlocks potential for scalable, efficient medical imaging solutions. Its architectural innovations—detection and context modules—combine local feature analysis with contextual understanding to produce interpretable, well-calibrated heatmaps. These heatmaps enhance diagnostic transparency and reduce reliance on pixel-level annotations. Such capabilities could significantly accelerate the annotation process (Raciti et al., 2020), allowing clinicians to refine rather than annotate from scratch, streamlining workflows requiring detailed labels.
@@
Limitations. While INSIGHT has been validated on CT and WSI data, its applicability to other modalities, such as MRI and ultrasound, remains unexplored. In principle, the architecture should generalize due to its ability to process spatially embedded features and generate interpretable outputs. However, a key barrier lies in the scarcity of public datasets. Future work will focus on identifying or curating suitable datasets to assess generalizability across broader modalities.
@@
Conclusion. INSIGHT represents a significant advancement in weakly supervised medical imaging, demonstrating strong performance across diverse tasks and modalities. By leveraging image-level labels, embedding interpretability directly into its architecture, and providing fine-grained diagnostic insights, INSIGHT lays the foundation for scalable, clinically relevant AI solutions. Future work will enhance its capabilities with advanced foundation models and extend its reach to additional modalities, broadening its impact on medical imaging research and practice.
[…]
@@
Appendix A Implementation Details
Architecture Parameters. INSIGHT uses UNI (Chen et al., 2024) and Virchow (Vorontsov et al., 2024) for WSI tiles, and ViT-L/DINOv2 (Oquab et al., 2024) for CT slices. Pre-processed inputs are resized to $224\times 224$, and embeddings from the last layer of the encoder are concatenated to retain spatial information. Dimensionality is reduced from $1024$ to $128$ using a linear transformation. Detection and Context modules consist of three convolutional layers with GELU activations, layer normalization, and a final layer for label-specific outputs. The Detection module uses $1\times 1$ convolutions, while the Context module uses $3\times 3$ convolutions.
Training Configuration. INSIGHT is trained using AdamW with a learning rate of $10^{-4}$ and weight decay of $10^{-4}$. Training is capped at $50$ epochs, with early stopping triggered after $8$ epochs of no validation improvement. We performed grid search to tune hyperparameters on the Camelyon16 validation set, including SmoothMax sharpness $\alpha$ (4, 6, 8, 10), learning rates (5e-5, 1e-4, 3e-4), and spectral decoupling strengths $\lambda$ (0, 0.01, 0.05). The selected configuration was then applied to BRACS and MosMed without further tuning, demonstrating INSIGHT’s robustness across tasks. To ensure fair comparison, all baseline models were implemented using the original configurations specified in their respective papers.
Appendix B Qualitative Analysis
@@
We provide additional visualization results (Fig. 6789) to further demonstrate INSIGHT’s ability to generate well-calibrated heatmaps that align with diagnostically relevant regions in the CAMELYON16 dataset. These examples include results using both UNI and Virchow 2 feature encoders, highlighting the consistency and generalizability of our method across different backbone models. Please zoom in to view finer details.
@@
Figure 6: Original histopathology image (left) and the corresponding heatmap generated by INSIGHT using the Virchow 2 encoder (right). The green outlines on the original image represent ground-truth regions. The bottom row provides zoomed-in views of the regions highlighted by yellow boxes.
@@
Figure 7: WSI heatmap visualization (right) generated by INSIGHT using Virchow 2 as the feature extractor on the CAMELYON16 official dataset. The green outlines on the original image (left) represent ground-truth regions.
Figure 8: WSI heatmap visualization (right) generated by INSIGHT using UNI as the feature extractor on the CAMELYON16 official dataset. The green outlines on the original image (left) represent ground-truth regions.
Figure 9: Additional WSI heatmap visualization (right) generated by INSIGHT using UNI as the feature extractor on the CAMELYON16 official dataset. The green outlines on the original image (left) represent ground-truth regions.
Appendix C Runtime Analysis
A key advantage of INSIGHT is native support for multi-label classification. In contrast, baselines like ABMIL and CLAM variants are inherently multi-class only and require running separate models per label for multi-label tasks such as BRACS. This makes INSIGHT significantly more efficient in real-world settings. We report runtime (min/epoch) for training and inference below.
Table 8: Runtime (min/epoch) comparison across datasets and models.
Dataset (task) Model Train Inference
Camelyon16 (binary) INSIGHT (Ours) 81.8 4.2
ABMIL 21.2 2.2
CLAM-SB 19.3 2.2
CLAM-MB 21.5 2.3
TransMIL 28.3 2.6
WiKG 24.7 2.4
BRACS (multilabel) INSIGHT (Ours) 114.7 6.8
ABMIL 111.5 12.8
CLAM-SB 112.8 15.3
CLAM-MB 121.4 16.6
TransMIL 132.7 19.2
WiKG 125.3 16.9
Appendix D Heatmap Generation Algorithm
Input: WSI tiles or CT slices $\{x_{i}\}_{i=1}^{N}$ with embeddings $\{e_{i}\}_{i=1}^{N}$, Detection module $f_{\text{Det}}$, Context module $f_{\text{Con}}$, and pooling sharpness $\alpha$.
Output: Prediction $\hat{y}$ and reconstructed heatmap $\mathcal{H}_{\text{full}}$.
Step 1: Feature Projection.
for $i=1$ to NN do
 $e_{i}\leftarrow\text{Conv}_{1\times 1}(e_{i})$
    // 1×1 projection to reduce dimension
end for
Step 2: Module Outputs.
for $i=1$ to NN do
 $\mathcal{H}_{\text{Det},i}\leftarrow f_{\text{Det}}(e_{i})$
    // Local detail map
 $\mathcal{H}_{\text{Con},i}\leftarrow f_{\text{Con}}(e_{i})$
    // Context map
 $\mathcal{H}_{i}\leftarrow\sigma\big{(}(1-\sigma(\mathcal{H}_{\text{Con},i}))\odot\mathcal{H}_{\text{Det},i}\big{)}$
end for
Step 3: Patch-to-Image Fusion.
Arrange $\{\mathcal{H}_{i}\}$ into a full heatmap $\mathcal{H}_{\text{full}}$ according to spatial coordinates.
Apply Otsu threshold $T\leftarrow\arg\min_{t}\sigma_{w}^{2}(t)$.
$\mathcal{H}_{\text{full}}\leftarrow\mathcal{H}_{\text{full}}\cdot\mathbb{I}(\mathcal{H}_{\text{full}}>T)$
Step 4: SmoothMax Pooling.
$\hat{y}\leftarrow\frac{\sum_{j}\mathcal{H}_{\text{full},j}\cdot\exp(\alpha\cdot\mathcal{H}_{\text{full},j})}{\sum_{j}\exp(\alpha\cdot\mathcal{H}_{\text{full},j})}$
Step 5: (Multi-label).
If multi-label, repeat Steps 2–4 for each class $c\in\{1,\dots,C\}$, where $C$ is the number of labels. This produces channel-wise outputs $\hat{y}_{c}$, resulting in a prediction vector $\hat{\mathbf{y}}\in\mathbb{R}^{C}$.
Algorithm 1 INSIGHT Heatmap Generation and Classification